OpenStudy (anonymous):

Sara makes and sells bracelets. She bought material for $28.50 and used it all to make 15 bracelets. Sara used the equation 15x-28.50=99 to determine x, the amount she should charge for each bracelet to make a profit of $99. How much should each bracelet cost? A. $4.70 B. $6.60 C. $8.50 D. $13.75

OpenStudy (anonymous):

c or b i think

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

you have the equation 15x-28.50=99 15x - 28.50 + 28.50 = 99 + 28.50 15x = 127.5 x = 25.5

OpenStudy (anonymous):

no not that

OpenStudy (anonymous):

x = 8.5

OpenStudy (anonymous):

which will be C

OpenStudy (anonymous):

thank you i thought it was B.

OpenStudy (anonymous):

Thank You Again!

OpenStudy (anonymous):

yw